Adults and high school students were asked four trivia questions. The number of correct answers given by each participant was recorded. To the right is a table showing the results. Complete parts​ a) through​ e).

,0,1,2,3,4,Total
Adults,16,32,36,45,31,160
Students,45,15,7,40,40,147
Total,61,47,43,85,71,307

a. If we select a participant at​ random, what's the probability we choose an adult who had one correct​ answer?

​b) Among the participants who had one correct​ answer, what is the probability the person is an​ adult?

​c)

What is the probability that an adult participant had one correct​ answer?

d. What is the probability that a student participant had no correct​ answers?

e. What is the probability that a participant with no correct answers is a​ student?

(Round to three decimal places as​ needed.)
Answer #1
0 1 2 3 4 total
adult 16 32 36 45 31 160
student 45 15 7 40 40 147
total 61 47 43 85 71 307

a)

probability we choose an adult who had one correct​ answer =32/307

b)

P(Adult |one correct answer)=32/47

c)

probability that an adult participant had one correct​ answer =32/160 =1/5

d)

probability that a student participant had no correct​ answers=45/147 =15/49

e)

probability that a participant with no correct answers is a​ student =45/61